mice pm report n.
Download
Skip this Video
Loading SlideShow in 5 Seconds..
MICE PM report PowerPoint Presentation
Download Presentation
MICE PM report

Loading in 2 Seconds...

play fullscreen
1 / 10

MICE PM report - PowerPoint PPT Presentation


  • 101 Views
  • Uploaded on

MICE PM report. Target News Outcome of Spectrometer Solenoid Review New plans for Steps II & III RF Power news. MICE CMPB, 4th December, 2009. Target News. Online target working well Now part of routine MICE shifts Regular inspections, nothing to report

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about 'MICE PM report' - arwen


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
mice pm report
MICE PM report
  • Target News
  • Outcome of Spectrometer Solenoid Review
  • New plans for Steps II & III
  • RF Power news

MICE CMPB, 4th December, 2009

target news
Target News
  • Online target working well
  • Now part of routine MICE shifts
  • Regular inspections, nothing to report
  • Offline target/re-engineering going OK
    • T3, all new parts are made except bearing liners
    • Waiting for plastic bearing material, Torlon/Vespel
    • Quite clear now that DLC coating of metallic bearings is problematic, so have dismissed it
    • Also DLC on DLC working in vacuum not such good idea
    • First thing is to re-build T2 with plastic bearings and run it and see......
    • Some minor open issues with plastics, but nothing thought serious
target news1
Target news
  • Plans and Dates (only based on informal discussion):
    • Mid January – aim to have T2 rebuilt with Plastic bearings
    • Mid January – March – Run T2 and then assemble T3 in identical format if worth doing
    • Easter shutdown – aim to have two identical targets to resume on-line/offline programme
  • Stator considerations:
    • In parallel to above, stator QA programme is underway:
      • Mechanical field survey rig (G.Barber) to assess magnetic assymetries
      • ASD at RAL have been asked to help Sheffield group with precise coil winding – square conductors and precision coil formers and uniform vacuum impregnation
  • Worry is that Jason Tarrant leaves MICE in January........
outcome of the spectrometer solenoid review
Outcome of the Spectrometer Solenoid Review
  • Very productive review held at Berkeley on 18th, 19th November, chaired by Pasquale Fabbricatori
  • Briefly the important points are:
    • The overall thermal balance of the magnet is still not completely understood, but this is being addressed by an ANSYS model and more experimental data
    • Given the circumstances and schedule constraints, the addition of a fourth single-stage CCR to the HTS lead area is the right thing to do
    • Operating margin of magnet two to be measured during cooldown
    • The more complete fix to magnet one should be continued with
spectrometer solenoid1
Spectrometer Solenoid
  • Dates and decision points:
    • December – cooldown and training of magnet #2
    • Mid January, 2010:
      • Magnet perfoms against acceptance criteria
      • Deliver to FNAL for field mapping, then RAL by mid April, or:
      • Fails to meet acceptance criteria, so:
      • More complete fix required, then approx ten months’ work
      • Likely outcome here is that magnets one & two deliveries converge
  • Important to remember: neither scenario will get Step II taking data before August 2010 (long ISIS shutdown)
  • So, the plan will have to change........
new plans for steps ii iii
New plans for Steps II & III
  • At MICE TB this Monday, agreed that STEP II is abandoned
  • And that we go straight to STEP III and have it ready for data-taking by February, 2011 (end of long shutdown)
  • The scientific case for Step III.1 (LiH solid absorber is re-visited)
  • Yet to be agreed by MICE EB – next week
  • FC #1 will arrive July ‘10, will need storage, but in reality the ‘spare’ time will be very useful
  • Actions:
    • Means we integrate two magnets in parallel
    • More space required, opened negotiations for R79
    • Installation staff profile will look different
    • Top-level schedule to be re-built around these dates
rf power
RF Power
  • Lifted from Andy Moss!
  • Retest of medium power amplifiers have been done in the last month
  • Electrical checks of large amplifier auxiliary electrical systems underway
  • CERN HT connection used to speed up testing
  • Connecting together medium and large power amplifiers
  • Should have results of testing with RF in December for entire system into test load
rf power1
RF Power

This is what we have to fit in R5.2,

More complex than it was, this is a huge unknown!

Really need to get on with design work as early as we can

rf power2
RF Power
  • Present RF project plan:
  • Test large 116 amplifier
  • Design hall coax in detail and work out how to fit it all in space for steps 5 and 6
  • Assemble CERN amplifier and test with current power supplies
  • Autumn 2010, move first amplifier system to MICE hall and install
  • Conclusion:
  • Large amplifier ready and will re-commission to 1MW using old tubes
  • Will run for as long as possible to develop understanding and look for weak points
  • CERN amplifiers on site awaiting build
  • Hall layout needs careful work
  • search for smaller key coax components if possible
  • Two new TH116 tubes have been delivered to DL, two more due soon